# ROSE: Revolutionizing Open-Set Dense Segmentation with Patch-Wise Perceptual Large Multimodal Model

### v2

@@ 3.1 Task Definition
Our goal is to propose a Segmentation LMM that can not only segment objects based on human prompts or predefined categories but also autonomously predict dense segmentation masks without any human-provided information. Therefore, we define a new task, Free-vocabulary Segmentation, where the model accepts task-level instructions (e.g., “Can you segment the foreground instance?”), and generates a dense segmentation $\hat{\mathbf{M}}\in\mathbb{R}^{N\times H\times W}$ along with the predicted category $\hat{\mathbf{y}}_{\text{txt}}$ in a generative manner.
@@ 3.2 Patch-wise Perception
Embedding-based mask generation, as proposed by LISA [38], offers a solution to mask prediction. However, directly stacking <SEG> tokens proves inadequate for dense object prediction (shown in Tab. 3). Contemporary LMMs typically rely on ViT-based encoders or raw image patches for feature processing, handling image data in a patch-wise manner. Inspired by the SOLO instance segmentation model [72], which divides images into grid-based predictions, our model takes image patches as fundamental prediction units, enabling object detection on a finer scale. Our patch-wise perception process is shown in Fig. 2a. Specifically, each patch predicts the following three components: 1) objectness score, indicates the probability of an object of interest being present within the current patch. 2) Mask embedding, serves as input to the SAM module for mask generation. 3) Category embedding, utilized for subsequent classification tasks. To implement this, we begin by dividing an input image $\mathbf{x}_{\text{img}}$ of size $L$$\times$$L$ into non-overlapping patches of size $p$$\times$$p$, yielding $S$$\times$$S$ patches, where $S=\lfloor\frac{L}{p}\rfloor$. Thus, the maximum number of predictions is $S^{2}$. If a target object’s center falls within the spatial region of a patch located at coordinates $(h,w)$, supervision is assigned to that patch and its surrounding patches. All the $S^{2}$ patches are first passed into vision encoder $\mathcal{F}_{\text{vis}}$, and the resulting visual features, combined with the task instruction $\mathbf{x}_{\text{txt}}$, are further processed by LLM $\mathcal{F}_{\text{llm}}$:
@@ 4.1 Experiment setting
Dataset and Task. Our training tasks and datasets are composed of the following: 1) Semantic segmentation, we use ADE20k [91], COCO-Stuff [6], and Mapillary [50] in this task. 2) Instance segmentation, In this task, we only use one dataset, COCO [44], in the training. 3) Referring segmentation, Following LISA, we use RefCLEF, RefCOCO, RefCOCO+ [35], and RefCOCOg [49]. 4) Segmentation Refinement. There are two batches of data: the first follows  [14] randomly added some holes and extra patches, and the second randomly shrinks or stretches the object area and keeps parts of this variation.
@@ A.2 Super-patch
A.3 Dataset and Task

### v3

@@ 3.1 Task Definition
Our goal is to propose a Segmentation LMM that can not only segment objects based on human prompts or predefined categories but also autonomously predict dense segmentation masks without any human-provided information. We refer this task as Free-vocabulary Segmentation, where the model accepts task-level instructions (e.g., “Can you segment the foreground instance?”), and generates a dense segmentation $\hat{\mathbf{M}}\in\mathbb{R}^{N\times H\times W}$ along with the predicted category $\hat{\mathbf{y}}_{\text{txt}}$ in a generative manner.
@@ 3.2 Patch-wise Perception
Embedding-based mask generation, as proposed by LISA [38], offers a solution to mask prediction. However, directly stacking <SEG> tokens proves inadequate for dense object prediction (shown in Tab. 3). Contemporary LMMs typically rely on ViT-based encoders or raw image patches for feature processing, handling image data in a patch-wise manner. Inspired by the SOLO instance segmentation model [72], which divides images into grid-based predictions, our model takes image patches as fundamental prediction units, enabling object detection on a finer scale. Our patch-wise perception process is shown in Fig. 2a. Specifically, each patch predicts the following three components: 1) objectness score, indicates the probability of an object of interest being present within the current patch. 2) Mask embedding, serves as input to the SAM module for mask generation. 3) Category embedding, utilized for subsequent classification tasks. To implement this, we begin by dividing an input image $\mathbf{x}_{\text{img}}$ of size $L$$\times$$L$ into non-overlapping patches of size $p$$\times$$p$, yielding $S$$\times$$S$ patches, where $S=\lfloor\frac{L}{p}\rfloor$. Thus, the maximum number of predictions is $S^{2}$. If a target object’s mass center falls within the spatial region of a patch located at coordinates $(h,w)$, supervision is assigned to that patch and all 8 patches surround it $(h\pm 1,w\pm 1)$. All the $S^{2}$ patches are first passed into vision encoder $\mathcal{F}_{\text{vis}}$, and the resulting visual features, combined with the task instruction $\mathbf{x}_{\text{txt}}$, are further processed by LLM $\mathcal{F}_{\text{llm}}$:
@@ 4.1 Experiment setting
Dataset and Task. Our training tasks and datasets are composed of the following: 1) Semantic segmentation, we use ADE20k [91], COCO-Stuff [6], and Mapillary [50] in this task. 2) Instance segmentation, In this task, we only use one dataset, COCO [44], in the training. 3) Referring segmentation, Following LISA, we use RefCLEF, RefCOCO, RefCOCO+ [35], and RefCOCOg [49]. 4) Segmentation Refinement. We generate defective masks from GT masks to collect training pairs in two ways: the first follows  [14] randomly added some holes and extra patches, and the second randomly shrinks or stretches the object area and keeps parts of this variation.
@@ A.2 Super-patch
A.3 Training convergence
We compared the loss convergence of ROSE with LISA. ROSE requires a bit more trainable parameters (4.8%) than LISA (3.7%), but it affects convergence little in the training stage according to Fig. 6.
Figure 6: Mask loss during training.
A.4 Dataset and Task
@@
Figure 7: Failure cases. We show some samples of typical failure scenario in the ADE20k dataset. Wrong classification results are labeled.
@@
Appendix B Visualizations
B.1 Failure cases
Fig. 7 shows the typical flaw caused by granularity differences (“plant” and “tree”). Due to the inherent limitations of existing evaluation methods. Despite the correctness, it is evaluated as wrong. We think it is an important point for future research.
B.2 Qualitative results